Grade 12IIT JEE Entrance Exam

A projectile is given an initial velocity of (i +2j) m/s, where i is along the ground and j is along the vertical. If g=10 m/s^2, the equation of its trajectory is:
a) y= 2x – 5x^2
b) y= x – 5x^2
c) 4y= 2x – 5x^2
d) y= 2x + 5x^2

Archit
9 Years agoGrade 12

1 Answer

Khimraj
ApprovedApproved Tutor Answer9 Years ago
x=1*ty= 2t-1/2gt²Now put x =t and g=10y = 2x-5x²Hope it clears. If still u have then please clarify. If u like answer then please approve the answer.